Your family plans to start a small business in your neighborhood. Your father borrows $15,000 from the bank at an annual interest rate of 7.25% rate for 36 months. What is the amount of interest he will pay on this loan?
Found 2 solutions by Adedeji Adenike, ikleyn:
Answer by Adedeji Adenike(1)

i=prt/100 p=15000,r=7.25%,t=36months I=15000*7.25*36/100 150*7.25*36=39,150

Answer by ikleyn(52832)

.

The post says nothing about compounding, so from the context we can assume, by default, that the loan is simple interest.


Then the interest  I = P*r*t, where


    P = $15000  is the loan amount,

    r = 7.25% = 0.0725 is the annual interest rate,

    t = 3 years  (36 months, time in the formula is in YEARS ).


So, according to the formula, the amount of interest is


    P = 15000*0.0725*3 = 3262.50 dollars.      ANSWER

Solved.

---------------

To see many other similar solved problems,  look into the lesson
    - Simple interest percentage problems
in this site.

Notice that the novice tutor,  Adedeji,  mistakenly used  months  when calculated the interest.
In this place,  the YEARS  should be used,  OBVIOUSLY.


/\/\/\/\/\/\/\/

Adedeji,  I saw and read your self-description.
Looking in your post,  I see that you really need help in  Math.

Which grade are you ?


Come again to the forum soon with your problems and questions - we,  the tutors,  will help you.


Do not hesitate to post your solutions to the forum --- it is very good way to develop your skills.